Find the slope of the line 8x + 4y = 9. Enter the exact answer as a decimal and show all your work.
Answer by stanbon(75887)

Find the slope of the line 8x + 4y = 9. Enter the exact answer as a decimal and show all your work.
----------------
Write the equation in slope-intercept form by solving for "y":
8x+4y = 9
Subtract 8x from both sides:
---
4y = -8x+9
Divide both sides by 4 to get:
-----
y = -2x+(9/4)
----
The slope-intercept form is y = mx + b
where m is the slope and b is the intercept.
----
slope = -2 or -2.00 in decimal form
=====================
Cheers,
Stan H.